OpenStudy (anonymous):

How can I simplify -p*log(p) - q*log(q) - (1-p-q)*log(1-p-q)

OpenStudy (anonymous):

where the logs are base 2

OpenStudy (tkhunny):

Not clear what would make it "simplified". Use some log rules and see if you like it. \(-p\log_{2}(p) = \log_{2}\left(p^{-p}\right) = \log_{2}\left(\dfrac{1}{p^{p}}\right)\) Again, it's not clear at all that this qualifies as "simplified".